find three consecutive odd integers such that the sum of the first, twice the second, and three times the third is 94
Answer by Alan3354(69443)

find three consecutive odd integers such that the sum of the first, twice the second, and three times the third is 94
---------------
The easy way:
The sum is aprox 6 times the 1st integer
94/6 = 15.xxx
--> 13, 15 & 17
------------------
The hard way:
n, n+2 and n+4 are the integers
n + 2(n+2) + 3(n+4) = 94
6n + 16 = 94
6n = 78
n = 13